If m varies directly as x and m = 60 when x = 5 then k, the constant of variation is..?

Answer by Ruli(21)

m varies directly as x = m=kx
m = 60 when x = 5 = 60=k(5)
60 = k(5)
Divide each side by 5
12 = k
The constant variation is 12
check:
60 = 12(5)
60 = 60